OpenStudy (anonymous):

Pleaseeeeee help Ben wishes to blend candy selling for $1.60 a pound with candy selling for $2.50 a pound to get a mixture that will be sold for 1.90 a pound. How many pounds of the $1.60 and the $2.50 candy should be used to get 30 pounds of the mixture?

OpenStudy (anonymous):

put \(x\) as the number of pounds of $2.50 candy, then since the total pounds is 30, the number of pounds of $1.60 candy is \(30-x\) the total cost will be \[2.50x+1.60(30-x)\] which you want to be \(1.90\times 30=57\) so set them equal \[2.50x+1.60(30-x)=57\] and solve for \(x\)

OpenStudy (anonymous):

probably easier to multiply by 10 and solve \[25x+16(30-x)=570\]